# Complexity Assessment of Analog and Digital Security Primitives Signals Using the Disentropy of Autocorrelation

### v2

@@ Abstract
The study of regularity in signals can be of great importance, typically in medicine to analyse electrocardiogram (ECG) or electromyography (EMG) signals, but also in climate studies, finance or security. In this work we focus on security primitives such as Physical Unclonable Functions (PUFs) or Pseudo-Random Number Generators (PRNGs). Such primitives must have a high level of complexity or entropy in their responses to guarantee enough security for their applications. There are several ways of assessing the complexity of their responses, especially in the binary domain. With the development of analog PUFs such as optical (photonic) PUFs, it would be useful to be able to assess their complexity in the analog domain when designing them, for example, before converting analog signals into binary. In this numerical study, we decided to explore the potential of the disentropy of autocorrelation as a measure of complexity for security primitives as PUFs, TRNGs or PRNGs with analog output or responses. We compare this metric to others used to assess regularities in analog signals such as Approximate Entropy (ApEn) and Fuzzy Entropy (FuzEn). We show that the disentropy of autocorrelation is able to differentiate between well-known PRNGs and non-optimised or bad PRNGs in the analog and binary domain with a better contrast than ApEn and FuzEn. Next, we show that the disentropy of autocorrelation is able to detect small patterns injected in PUFs responses and then we applied it to photonic PUFs simulations.
@@ I Introduction
There is an entire zoo of PUFs, most of which are electronic ones and operate in the binary domain as the Arbiter, SRAM, or Butterfly PUFs [7]. Most of these electronic PUFs have responses of one bit for one given challenge. A multi-bits sequence is therefore built by concatenating the one-bit responses of the PUF for different challenges. In that case there is no single unambiguous way to assemble its responses into a sequence. For these PUFs, the NIST test suite [8], especially suited for the analysis of binary data, can be used to evaluate some of the statistical properties of the responses [9]. However, PUFs can also operate in the analog domain, especially in some recent implementations of optical and photonic PUFs [3]. For some of these PUFs one challenge give a ordered response of multiple bits [10, 11, 12, 13, 14].
Nevertheless, there is an important question to address when considering these PUFs that operate in the analog domain. The impact of fabrication variations affect some physical parameters, but PUFs output signals may present the need for a conversion from the analog to the binary domain to obtain responses that can be further used at a system level [10, 11, 12, 13, 14]. The quality of such responses will be impacted by the chosen conversion scheme: a different scheme will lead to different responses and probably different responses quality. Here by responses quality we refer to the responses complexity and entropy. Besides, post-processing may be needed to improve the responses quality. For example, in the first optical PUF by Pappu and co-workers in 2002 [13], the optical output of the PUF is a speckle pattern recorded on a $320\times 240$ pixel camera that passes through a threshold filter to obtain an output in the binary domain, and finally through a Gabor transform to obtain a 2400-bit key. As discussed in [14], the entropy of the output response strongly depends on this transformation. A Gabor transformed image has regularities (zebra-stripes) leading to important patterns in the responses and hence to low entropy. Therefore, the Gabor transform may remove parts of the complexity contained in the optical speckle pattern.
@@ V-A Tests on analog signals
We observe on Fig. 1a that the output of MT0 does not exhibit observable patterns. On the other hand, by looking at Fig. 1b we see that LCG Bad has clear repeating patters and should have bad results for all the metrics, i.e, ApEn and FuzEn close to zero and a high disentropy. As seen on Fig. 1c, LCG 2 does not necessarily have repeating patterns, but straight lines can be observed in its output.
@@ V-B Tests on binary signals
This algorithm obtains results for $m=2,3$ and $8$ similar to the MATLAB™
approximateEntropy function. In terms of $p$-values, only LCG Bad is considered as non-random except for $m=8$ where LCG 1 obtains $p=0.0096$ and is therefore near the limit.
@@ VI-A Deterministic patterns
VI-B Test on PUFs responses
Here, we use the simulated outputs of photonic PUFs previously published [11]. In this paper different photonic PUF architectures based on symmetric ring resonator arrays have been simulated and their responses have been tested using NIST tests (including ApEn) for two different binary conversions. We will focus our discussion here on PUF 1 and PUF 3 because of their different NIST ApEn [48] scores. PUF 3 is very simple and composed of one ring per array while PUF 1 is more complex and composed of 10 rings per array.
In this paper, the binary challenges are created using the MT algorithm and encoded in light. 150 instances of each architecture are simulated with fabrication variations. For each instance 100 responses of 128 bits are generated by sampling and converting the analog outputs into binary. At the end, these responses have been concatenated for the NIST tests evaluation. Their results are compared in Table VI with $N_{ApEn}^{\%}$ the percentage of instances passing the ApEn test [48].
TABLE VI: Results summary of PUF 1 and PUF 3 architectures.
PUF 1 PUF 3
$N_{ApEn}^{\%}$ 1-7% 46%
Uniqueness High Low
Impact on challenge-encoded inputs High Low
For two different binary conversion schemes, PUF 1 obtains low $N_{ApEn}^{\%}$ compared to PUF 3 despite the bias in PUF 3 responses due to the conversion to a binary format. PUF 1 can operate effectively on the inputs encoded with the challenges to generate unique instances because of its fabrication variations and architecture. It also degrades the quality of the inputs, while PUF 3 is not complex enough to have an impact on the inputs.
[…]
Also, as shown in Fig. 3(a), the score $\mathcal{D}(\textnormal{PUF 1})\sim 10^{-2}$ has the same order of magnitude as $\mathcal{D}$ of LCG 1, 3 and 4 exhibiting patterns. This indicates the presence of patterns in the analog PUF 1 outputs and indicates that the low $N_{ApEn}^{\%}$(PUF 1) is not entirely caused by a bad binary conversion scheme or the bias in the responses.
On the ApEn and FuzEn side, we discovered that the analog outputs of PUF 3 obtained a score very close to $\log(2)$ for both metrics. This means that ApEn and FuzEn are seeing the outputs of PUF 3 as signals close to binary data.
As shown in Fig. 6, the normalised analog outputs of PUF 3 are closely oscillating between two levels: 0 and 1 due to the low impact that the PUF 3 architecture has on the input light encoded with the binary challenges. Therefore, ApEn(PUF 1) $>$ ApEn(PUF 3) and FuzEn(PUF 1) $>$ FuzEn(PUF 3) even if PUF 1 has patterns and $N_{ApEn}^{\%}$(PUF 1) $<$ $N_{ApEn}^{\%}$(PUF 3).
Figure 6: Normalised output 1 of PUF 3 first instance compared to challenge 1. The delay caused by the PUF have been compensated to align the challenge with the output on the figure.
Hence, one should be careful while using ApEn and FuzEn on analog signals if they oscillate between two levels. As we observed on Fig 3(a) and Fig. 5(a), the disentropy is not highly affected by the transformation from analog to binary signal (c.f the well-known PRNGs and LCG Bad results). As mentioned in [32], the ApEn algorithm gives a relative value allowing to compare signals using the same “alphabet”. Therefore, depending on the application, this characteristic of ApEn and FuzEn to be sensitive to the type of signals can make a comparison between analog signals complicated.
One could argue that a binary signal is less complex than an analog signal with a very large number of levels. However, for security applications where the signal has to be converted into binary data, the detection of patterns and correlations in the signal might be more important than the complexity in terms of levels and the instability of ApEn and FuzEn can be detrimental in this regard.
@@ VII Conclusion
However, we saw that none of the metrics were able to detect the lines in LCG 2 analog output. ApEn and FuzEn failed in detecting these lines for a period $p_{line}>m$ and the disentropy for $p_{line}>40$. Also, the score given by the disentropy depends highly on the length of the signal. Therefore, we recommend using it with a comparison. We suggest to compare the score obtained by the signal under test alongside a signal of the same length generated with the MT0 algorithm.
@@ Appendix B LCG Bad scores
The scores obtained by LCG Bad are hard to represent in Fig. 3 and Fig. 5 without losing visibility for the other PRNGs. Therefore, its results are presented in Tab. VII for analog signals and in Tab. VIII for binary signals.
TABLE VII: LCG Bad scores for analog signals.
@@
TABLE VIII: LCG Bad scores for binary signals.

### v3

@@ Abstract
The study of regularity in signals can be of great importance, typically in medicine to analyse electrocardiogram (ECG) or electromyography (EMG) signals, but also in climate studies, finance or security. In this work we focus on security primitives such as Physical Unclonable Functions (PUFs) or Pseudo-Random Number Generators (PRNGs). Such primitives must have a high level of complexity or entropy in their responses to guarantee enough security for their applications. There are several ways of assessing the complexity of their responses, especially in the binary domain. With the development of analog PUFs such as optical (photonic) PUFs, it would be useful to be able to assess their complexity in the analog domain when designing them, for example, before converting analog signals into binary. In this numerical study, we decided to explore the potential of the disentropy of autocorrelation as a measure of complexity for security primitives as PUFs, TRNGs or PRNGs with analog output or responses. We compare this metric to others used to assess regularities in analog signals such as Approximate Entropy (ApEn) and Fuzzy Entropy (FuzEn). We show that the disentropy of autocorrelation is able to differentiate between well-known PRNGs and non-optimised or bad PRNGs in the analog and binary domain with a better contrast than ApEn and FuzEn. Then, we show that the disentropy of autocorrelation is able to detect small patterns injected in PUFs responses.
@@ I Introduction
There is an entire zoo of PUFs, most of which are electronic ones and operate in the binary domain as the Arbiter, SRAM, or Butterfly PUFs [7]. Most of these electronic PUFs have responses of one bit for one given challenge. A multi-bits sequence is therefore built by concatenating the one-bit responses of the PUF for different challenges. In that case there is no single unambiguous way to assemble its responses into a sequence. For these PUFs, the NIST test suite [8], especially suited for the analysis of binary data, can be used to evaluate some of the statistical properties of the responses [9]. However, PUFs can also operate in the analog domain, especially in some recent implementations of optical and photonic PUFs [3]. For some of these PUFs one challenge give a ordered response of multiple bits [10, 11, 12, 13].
Nevertheless, there is an important question to address when considering these PUFs that operate in the analog domain. The impact of fabrication variations affect some physical parameters, but PUFs output signals may present the need for a conversion from the analog to the binary domain to obtain responses that can be further used at a system level [10, 11, 12, 13]. The quality of such responses will be impacted by the chosen conversion scheme: a different scheme will lead to different responses and probably different responses quality. Here by responses quality we refer to the responses complexity and entropy. Besides, post-processing may be needed to improve the responses quality. For example, in the first optical PUF by Pappu and co-workers in 2002 [12], the optical output of the PUF is a speckle pattern recorded on a $320\times 240$ pixel camera that passes through a threshold filter to obtain an output in the binary domain, and finally through a Gabor transform to obtain a 2400-bit key. As discussed in [13], the entropy of the output response strongly depends on this transformation. A Gabor transformed image has regularities (zebra-stripes) leading to important patterns in the responses and hence to low entropy. Therefore, the Gabor transform may remove parts of the complexity contained in the optical speckle pattern.
@@ V-A Tests on analog signals
We observe on Fig. 1a that the output of MT0 does not exhibit observable patterns. On the other hand, by looking at Fig. 1b we see that LCG Bad has clear repeating patterns and should have bad results for all the metrics, i.e, ApEn and FuzEn close to zero and a high disentropy. As seen on Fig. 1c, LCG 2 does not necessarily have repeating patterns, but straight lines can be observed in its output.
@@
The metrics discussed thus far can also be applied to TRNGs allowing us to compare with other generators. Here we used a TRNG based on atmospheric noise [47]. As presented in Appendix C, for the different metrics, the TRNG obtains results similar to MT0 and MTS. This means that by taking only 10000 samples, much smaller than the period of MT, the different metrics are not able to detect patterns in MT outputs and consider it similar to a TRNG.
@@ V-B Tests on binary signals
This algorithm obtains results for $m=2,3$ and $8$ similar to the MATLAB™ approximateEntropy function. In terms of $p$-values, only LCG Bad is considered as non-random except for $m=8$ where LCG 1 obtains $p=0.0096$ and is therefore near the limit.
@@
V-C Test on multi-level analog signals
In this section we want to investigate how the different metrics are impacted by the number of levels, as we saw previously, ApEn and FuzEn have their results going from approximately $2$ for MT0 analog signal and $\log(2)$ for binary signals. Here, we will treat the different signals as if they were analog signal oscillating between multiple levels. To do so, we generate an analog signal of 10000 samples with MT0 and cut the interval [0,1] into multiple levels. The numbers generated with MT0 are assigned to a level depending on their value. The analysis is performed from two levels $\{0,1\}$ to ten levels $\{0,0.1111,0.2222,0.3333,0.4444,0.5556,0.6667,0.7778,\\ 0.8889,1\}$. Furthermore in order to take into account the fact that an analog signal can oscillate around multiple levels, we add a small Gaussian noise (MT) around the levels using the MATLAB™ normrnd function with $\mu=0$ and a $\sigma\in[0,0.1]$.
Figure 6: Evolution of a) ApEn$(m=3)$ b) FuzEn$(m=3)$ c) $\mathcal{D}$ with the number of levels of a multi-level signal generated MT0. A Gaussian noise with standard deviation $\sigma$ has been added to each level.
For ApEn, FuzEn we used $m=3$. As shown on Fig. 6a) and b), ApEn and FuzEn start around $\log(2)$ with two levels and $\sigma=0$, ApEn reaches a plateau around $1.8$ and FuzEn $1.9$ for ten levels giving results close to analog MT0 results of Tab.II. The fact of adding noise around the levels make ApEn and FuzEn perceive the signal “more analog” for ApEn and FuzEn. ApEn and FuzEn do not seem to be the metric to use when the frontier between analog and binary is less explicit than in the case where we clearly have ’0’ and ’1’ and in the case we have an analog signal oscillating between a significant amount of levels. On the contrary and as already observed on Fig 3(a) and Fig. 5(a), the disentropy is not highly affected by the transformation from analog to binary signal (c.f the well-known PRNGs and LCG Bad results). It is not very affected by the noise neither. As the multi-level signal is generated using MT0, therefore with the same amount of regularities as MT0, we expect a disentropy score close to the one of MT0. The fact of adding multiple sub-levels with the Gaussian noise using MT does not affect the disentropy neither.
Hence, one should be careful while using ApEn and FuzEn on analog signals if they oscillate between two (few) levels. As mentioned in [31], the ApEn algorithm gives a relative value allowing to compare signals using the same “alphabet”. Therefore, depending on the application, this characteristic of ApEn and FuzEn to be sensitive to the type of signals can make a comparison between analog signals complicated. Furthermore, as seen on Tab. 7 and Fig. 5 the choice of $m$ depends on the type of signal (analog/binary) and tests should be performed to find the correct $m$ to apply on the signal.
One could argue that a binary signal is less complex than an analog signal with a very large number of levels. However, for security applications where the signal has to be converted into binary data, the detection of patterns and correlations in the signal may be more important than the complexity in terms of levels and the instability of ApEn and FuzEn can be detrimental in this regard.
@@ VII Conclusion
However, we saw that none of the metrics were able to detect the lines in LCG 2 analog output. ApEn and FuzEn failed in detecting these lines for a period $p_{line}>m$ and the disentropy for $p_{line}>40$. Also, the score given by the disentropy depends highly on the length of the signal. Therefore, we recommend using it with a comparison. We suggest to compare the score obtained by the signal under test alongside a signal of the same length generated with the MT0 algorithm or a TRNG like an atmospheric noise-based TRNG [47].
@@ Appendix B LCG Bad scores
The scores obtained by LCG Bad are hard to represent in Fig. 3 and Fig. 5 without losing visibility for the other PRNGs. Therefore, its results are presented in Tab. VI for analog signals and in Tab. VII for binary signals.
TABLE VI: LCG Bad scores for analog signals.
@@
TABLE VII: LCG Bad scores for binary signals.
@@
Appendix C Atmospheric noise TRNG
In this section we will generate sequences from data generated from the website at Ref. [47]. It uses a TRNG based on atmospheric noise. The service has been built by Dr. Mads Haahr of the School of Computer Science and Statistics at Trinity College in Dublin and is now operated by Randomness and Integrity Services Ltd [47]. Atmospheric noise is generated by lightning discharges. A lightning is not composed of one plasma pulse but is a complex phenomenon composed of multiple steps and pulses generating the atmospheric noise [49, 50].
Since we cannot set a seed and collect the same sequence for this TRNG it is necessary to collect multiple sequences and test how they perform on average for multiple sequences generated with the TRNG. In order to compare this result with the PRNGs we decided to generate sequences of 10000 samples using the “Random Integer Set Generator”. Each of these sets contains integers ranging from 1 to 1,000,000. The sets are collected and normalized in order to obtain a signal with 10000 samples $\in[0,1]$ before testing them with the metrics. The number of sets/samples is limited on the website, therefore we decided to limit this statistical study to 10 sets. The reader may argue that this could be too low but it would already give a good idea of the TRNG performance. Since we want to detect patterns with 10000 samples we expect the results of the TRNG to be similar to the ones of MT0 and MTS presented in Tab. II and Fig. 3. TRNG results are presented on Tab. VIII.
TABLE VIII: TRNG score
$\mathcal{D}$ ApEn FuzEn
$(2.10\pm 1.40)\cdot 10^{-4}$ $m=2$ $2.160\pm 0.003$ $2.042\pm 0.003$
$m=3$ $1.846\pm 0.008$ $1.923\pm 0.004$
These results are indeed similar to those obtained by the MT0 and MT0 algorithms. Therefore, we can conclude that as long as the sequence length is much smaller than the maximum period of MT, the metrics consider that MT generates a random sequences.